40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1618     Entry time: Thu May 21 18:21:57 2009
Author: rana 
Type: Summary 
Category: Treasure 
Subject: Yoichi's words 

Yoichi's final words on what do next with the interferometer (as of 5 PM on May 21, 2009):

  1. Measure laser noise couplings in spring and anti-spring configurations.
  2. Dewhitening filter turn on for the ETMs.
  3. Noisebudget - import from the sites.
  4. Stabilize CM handoff.

My personal sub-comments to these bullets:

  1. For the laser noise I'm not sure that we will be able to understand these if the couplings are mainly from junk light due to accidental HOM resonances.
  2. WE should look into putting a static passive stage of filtering into the ETMs if warranted by the NB.
  3. Because of the sad track record with this, I will start us off this time by importing and modifying the H1/L1 versions.
  4. I guess we can do this by just acquiring on MC2 with the huge CARM offset. It works for the single arm so it should work for offset CARM.
ELOG V3.1.3-